Enviro Company issues 8%, 10-year bonds with a par value of $260,000 and semiannual interest payments. On the issue date, the annual market rate for these bonds is 10%, which implies a selling price of 87 1⁄2. The straight-line method is used to allocate interest expense.

1. Using the implied selling price of 87 ½, what are the issuer's cash proceeds from issuance of these bonds?
2. What total amount of bond interest expense will be recognized over the life of these bonds?
3. What is the amount of bond interest expense recorded on the first interest payment date?
Answer #1

1) Net Cash Proceeds = 260,000*0.875 = $ 227,500 (Answer)

2). Total bond interest Expenses

Cash interest = 260,000*8%*10 years =208,000

Discount on Bond Issue = 260,000-227,500 = $32,500

Total bond Interest Expneses = 208,000+32,500 = $ 240,500 (Answer)

3) Interest Expneses on First interest Paymnet :

Cash interest = 260,000*8%*6/12 =10,400

Discount Amotized = 32,500/20 =1,625

Interest Expenses = $12,025 (answer)

Face value of the Bond Less: Bond Selling price (260000*0.875) Dicount on bonds Payable 260000 227500 32500 Cash Proceeds 227